# NeuronSeek: On Stability and Expressivity of Task-driven Neurons

### v1

@@ V-D Comparison of Tensor Decomposition Methods
Our framework employs the CP decomposition for neuronal discovery. To empirically validate this choice, we conduct a comparative analysis of some tensor decomposition methods used in NeuronSeek-TD.
V-D1 Experimental settings
@@
Generally, we follow the previous experimental settings in Section V-C: 16 tabular datasets and the same network structures as described in Table III. We adopt CP decomposition and Tucker decomposition respectively for tensor processing in NeuronSeek-TD. We employ MSE for regression tasks, while classification accuracy is utilized for classification tasks.
Figure 9: Comparison between CP and Tucker decomposition. (a) Performance comparison on regression tasks measured by MSE. (b) Performance comparison on classification tasks measured by accuracy.
Figure 10: Comparison between CP and Tucker decomposition
V-D2 Experimental results
@@
First, Figure 9 illustrates that, for most datasets, the two decomposition methods exhibit comparable performance. Exceptions include airfoil Self-Noise and electricity datasets, where the two methods show slight performance differences. This observation suggests that variations in TD methods do not substantially impact the final results. Second, as shown in Figure 10, the CP-based method achieves a 42% reduction in computational time. This result verifies that the CP method offers a computational advantage over the Tucker method.
VI Competitive Experiments of NeuronSeek-TD over Other Baselines
VI-A Experiment on Tabular Data
@@
To further illustrate the superiority of NeuronSeek-TD, we compare it with other state-of-the-art models over 4 real-world tabular tasks: Electron Collision Prediction [^1] https://www.kaggle.com/datasets/fedesoriano/cern-electron-collision-data, Asteroid Prediction [^2] https://www.kaggle.com/datasets/basu369victor/prediction-of-asteroid-diameter, Heart Failure Detection [^3] https://www.kaggle.com/datasets/fedesoriano/heart-failure-prediction/data, and Stellar Classification [^4] https://www.kaggle.com/datasets/fedesoriano/stellar-classification-dataset-sdss17). We adopt 7 advanced machine learning methods, namely XGBoost [59], LightGBM [60], CatBoost [61], TabNet [62], TabTransformer [63], FT-Transformer [64], DANETs [65] and NeuronSeek-SR [12] as the baselines to compare with NeuronSeek-TD. Moreover, our proposed model and NeuronSeek-SR replace standard neurons with what are discovered by NeuronSeek-TD and NeuronSeek-SR, respectively.
@@
All datasets are randomly split into training, validation, and test sets in a 0.8:0.1:0.1 ratio. For regression tasks (electron collision and asteroid prediction), we use the MSE as an evaluation metric, while for classification tasks with imbalanced labels (heart failure detection and stellar classification), we adopt the F1 score. Detailed experiment information is shown in Table IV.
TABLE IV: Details of real-world tabular datasets.
@@
TABLE V: The test results of different models on real-world datasets.
@@
NeuronSeek-SR 0.0016 $\pm$ 0.0005 0.0513 $\pm$ 0.0551 0.8874 $\pm$ 0.05 0.9613 $\pm$ 0.002
NeuronSeek-TD (ours) 0.0011 $\pm$ 0.0003 0.0502 $\pm$ 0.0800 0.9023 $\pm$ 0.03 0.9714 $\pm$ 0.001
Results are compared in Table V. It can be seen that TabTransformer demonstrates competitive performance on the particle collision dataset. However, its effectiveness sharply declines on the asteroid prediction. In contrast, CatBoost and TabNet exhibit more consistent yet suboptimal results across four datasets. Lastly, the proposed NeuronSeek-TD method is the best performer, achieving the lowest error rates and highest F1 scores uniformly across all four benchmarks. Notably, NeuronSeek-TD significantly outperforms FT-Transformer and DANET. This significant margin underscores the efficacy of its task-adaptive architecture in recognizing feature interactions without domain-specific tuning.
VI-B Experiments on Images
To evaluate the performance enhancement of NeuronSeek-TD in computer vision tasks, we integrate it into several established convolutional neural network architectures: ResNet [3], DenseNet [66], SeResNet [67], and GoogLeNet [68]. Notably, we replace the first convolutional layer of each network block with a TN convolutional layer employing the discovered formula $x+\sin x$. It should be noted that the original NeuronSeek-SR was not designed for high-dimensional image tasks due to convergence difficulties. Our implementation focuses on NeuronSeek-TD and follows the standardized training protocol from a popular image classification repository [^5] https://github.com/weiaicunzai/pytorch-cifar100, incorporating several advanced training techniques: Learning rate scheduling via WarmUpLR and MultistepLR, regularization through Label Smoothing, and optimization using SGD. We use grid search to fine-tune the learning rate from 0.5 to 0.05. All models are trained for 200 epochs with a batch size of 128. We evaluate the results using classification accuracy averaged over 10 independent runs.
We adopt four standard image datasets to validate the model: CIFAR 10/100 [69], SVHN [70], and STL10 [71]. The training and validation sets are split with a ratio of 0.9 : 0.1. In particular, images from the STL10 dataset are resized to $32\times 32$ to fit the input of the model. The classification results are summarized in Table VI.
TABLE VI: Accuracy of compared methods on small-scale image datasets.
@@ VI-B Experiments on Images
VII Discussions and Conclusions
In this paper, we have proposed a novel method in the framework of NeuronSeek to enhance the capabilities of a neural network. Compared to neural architecture search (NAS) [72], NeuronSeek provides a different approach. Rather than modifying the overall model structure, NeuronSeek optimizes solely at the neuronal level by searching for optimal aggregation formulas. Firstly, we have introduced tensor decomposition to prototype stable task-driven neurons. Secondly, we have provided theoretical guarantees demonstrating that these neurons can maintain the super-super-expressive property. At last, through extensive empirical evaluation, we show that our proposed framework consistently outperforms baseline models while achieving the state-of-the-art results across diverse datasets and tasks. These advancements establish task-driven neurons as an effective addition in the design of task-specific neural network components, offering both theoretical soundness and practical efficacy.

### v2

@@ 5.4 Comparison of Tensor Decomposition Methods
Our framework employs the CP decomposition for neuronal discovery. To empirically validate this choice, we conduct a comparative analysis of some tensor decomposition methods used in NS-TD.
5.4.1 Experimental Settings
@@
Generally, we follow the previous experimental settings in Section 5.3: 16 tabular datasets and the same network structures as described in Table 5. We adopt CP decomposition and Tucker decomposition, respectively, for tensor processing in NS-TD. We employ MSE for regression tasks and accuracy for classification tasks.
5.4.2 Experimental Results
@@
First, Fig. 10 illustrates that, for most datasets, the two decomposition methods exhibit comparable performance. Exceptions include airfoil Self-Noise and electricity datasets, where two methods show slight performance differences. This observation suggests that variations in decompositions for TD methods do not substantially impact the final results. Second, as shown in Fig. 11, the CP-based method achieves a 42% reduction in computational time. This result verifies that the CP method offers a computational advantage over the Tucker method.
(a) Regression Task
(b) Classification Task
Figure 10: Comparison between CP and Tucker decomposition.
Figure 11: Time comparison between CP and Tucker decomposition.
6 Competitive Experiments of NS-TD over Other Baselines
6.1 Experiment on Tabular Data
@@
To further illustrate the superiority of NS-TD, we compare it with other state-of-the-art models over 4 real-world tabular tasks: Electron Collision Prediction, Asteroid Prediction, Heart Failure Detection, and Stellar Classification). We adopt 7 advanced machine learning methods, namely XGBoost [6], LightGBM [30], CatBoost [21], TabNet [1], TabTransformer [26], FT-Transformer [19], DANETs [5] and NS-SR [17] as the baselines to compare with NS-TD. Moreover, our proposed model and NS-SR replace standard neurons with what are discovered by NS-TD and NS-SR, respectively.
@@
All datasets are randomly split into training, validation, and test sets in a 0.8:0.1:0.1 ratio. For regression tasks (electron collision and asteroid prediction), we use the MSE as an evaluation metric, while for classification tasks with imbalanced labels (heart failure detection and stellar classification), we adopt the F1 score. Detailed experiment information is shown in Table 6.
Table 6: Details of real-world tabular datasets.
@@
In Table 7, it can be seen that TabTransformer demonstrates competitive performance on the electron collision dataset. However, its effectiveness sharply declines on the asteroid prediction. In contrast, CatBoost and TabNet exhibit more consistent yet suboptimal results across four datasets. Lastly, the proposed NS-TD is the best performer, achieving the lowest error rates and highest F1 scores uniformly across all four benchmarks. Notably, NS-TD significantly outperforms FT-Transformer and DANET. This significant margin underscores the efficacy of its task-adaptive architecture in recognizing feature interactions without domain-specific tuning.
Table 7: The test results of different models on real-world datasets.
@@
NS-SR 0.0016 $\pm$ 0.0005 0.0513 $\pm$ 0.0551 0.8874 $\pm$ 0.05 0.9613 $\pm$ 0.002
NS-TD (ours) 0.0011 $\pm$ 0.0003 0.0502 $\pm$ 0.0800 0.9023 $\pm$ 0.03 0.9714 $\pm$ 0.001
6.2 Experiments on Images
To evaluate the performance enhancement of NS-TD in computer vision tasks, we integrate it into several established convolutional neural network architectures: ResNet [22], DenseNet [24], SeResNet [23], and GoogLeNet [56]. Notably, we replace the first convolutional layer of each network block with a TN convolutional layer employing the discovered formula $x+\sin x$. It should be noted that the original NS-SR was not designed for high-dimensional image tasks due to convergence difficulties. Our implementation focuses on NS-TD and follows the standardized training protocol from a popular image classification repository [^3] https://github.com/weiaicunzai/pytorch-cifar100, incorporating several advanced training techniques: Learning rate scheduling via WarmUpLR and MultistepLR, regularization through Label Smoothing, and optimization using SGD. We use grid search to fine-tune the learning rate from 0.5 to 0.05. All models are trained for 200 epochs with a batch size of 128. We evaluate the results using classification accuracy averaged over 10 independent runs.
We adopt four standard image datasets to validate the model: CIFAR 10/100 [34], SVHN [46], and STL10 [11]. The training and validation sets are split with a ratio of 0.9:0.1. In particular, images from the STL10 dataset are resized to $32\times 32$ to fit the input of the model. The classification results are summarized in Table 8.
Table 8: Accuracy of compared methods on small-scale image datasets.
@@ 6.2 Experiments on Images
7 Conclusions
In this paper, we have proposed a novel method in the framework of NeuronSeek to enhance the capabilities of a neural network. Compared to neural architecture search (NAS) [61], NeuronSeek emphasizes the importance of neurons by searching for optimal aggregation formulas. Firstly, we have introduced tensor decomposition to prototype stable task-driven neurons. Secondly, we have provided theoretical guarantees demonstrating that these neurons can maintain the super-super-expressive property. At last, through extensive empirical evaluation, we have shown that our proposed framework consistently outperforms baseline models while achieving the state-of-the-art results across diverse datasets and tasks. These advancements establish task-driven neurons as an effective addition in the design of task-specific neural network components, offering both theoretical soundness and practical efficacy and stability.
Appendix A Supplementary Results
A.1 Experiments on Synthetic Data
Supplementary results on synthetic data (Section 5.1) are illustrated as follows.
Table 9: The reconstruction accuracy ($\text{Mean}_{\pm\text{SEM}}$) of pure mode.
Method Pure Mode ($\sum x^{k}$)
10 30 50 100
SR (PySR) $1612_{0.0224} 1.1$$1493_{0.0061}$ $1.2325_{0.0042}$ $1.2481_{0.0093}$
EQL $0.4524_{0.0232}$ $0.5342_{0.0141}$ $0.5705_{0.0123}$ $0.6923_{0.0132}$
NS-SR (VSR) $0.2324_{0.0041}$ $0.2541_{0.0092}$ $0.2825_{0.0063}$ $0.2953_{0.0071}$
MetaSymNet $0.8324_{0.0642}$ $0.8541_{0.0053}$ $0.9825_{0.0081}$ $0.9843_{0.0092}$
NS-TD $\mathbf{0.0543}_{0.0062}$ $\mathbf{0.0632}_{0.0051}$ $\mathbf{0.0854}_{0.0083}$ $\mathbf{0.0891}_{0.0082}$
Table 10: The reconstruction accuracy ($\text{Mean}_{\pm\text{SEM}}$) of interact mode.
Method Interact Mode (Vector Interact)
10 30 50 100
SR (PySR) $0.8754_{0.0102}$ $1.0042_{0.0084}$ $0.9776_{0.0192}$ $0.9763_{0.0091}$
EQL $0.3875_{0.0174}$ $0.4862_{0.0021}$ $0.5614_{0.0042}$ $0.4521_{0.0073}$
NS-SR (VSR) $0.6972_{0.0053}$ $0.7734_{0.0052}$ $0.7326_{0.0074}$ $0.8145_{0.0042}$
MetaSymNet $0.8632_{0.0064}$ $1.0804_{0.0093}$ $1.0193_{0.0102}$ $0.9942_{0.0073}$
NS-TD $\mathbf{0.0512}_{0.0041}$ $\mathbf{0.0583}_{0.0082}$ $\mathbf{0.0724}_{0.0053}$ $\mathbf{0.0945}_{0.0042}$
Table 11: The reconstruction accuracy ($\text{Mean}_{\pm\text{SEM}}$) of hybrid mode.
Method Hybrid Mode (Mixture)
10 30 50 100
SR (PySR) $1.0635_{0.0083}$ $0.9821_{0.0072}$ $0764_{0.0123} 1.1$$1018_{0.0241}$
EQL $0.3126_{0.0052}$ $0.4113_{0.0051}$ $0.4237_{0.0063}$ $0.4472_{0.0061}$
NS-SR (VSR) $0.7783_{0.0081}$ $0.7824_{0.0072}$ $0.8281_{0.0031}$ $0.8685_{0.0083}$
MetaSymNet $1.1345_{0.0124}$ $0.9526_{0.0082}$ $1093_{0.0173} 1.1$$1334_{0.0162}$
NS-TD $\mathbf{0.0423}_{0.0041}$ $\mathbf{0.0621}_{0.0043}$ $\mathbf{0.0735}_{0.0071}$ $\mathbf{0.1112}_{0.0012}$
A.2 Experiments on Tabular Data
We present the supplementary results of Section 5.3 to showcase the superiority of the proposed NS-TD.
Table 12: The supplementary results with standard deviations on the tabular datasets. The metrics are presented in the format $\text{Mean}_{\text{Std}}$.
Dataset MLP KAN NS-SR NS-TD (Ours)
Regression Tasks (Metric: MSE $\downarrow$)
California Housing $\mathbf{0.0890}_{0.0021}$ $0.1017_{0.0035}$ $0.1135_{0.0052}$ $\underline{0.1005}_{0.0050}$
[…]
Airlines Delay $0.1605_{0.0050}$ $\underline{0.1563}_{0.0055}$ $0.1602_{0.0055}$ $\mathbf{0.1560}_{0.0052}$
Classification Tasks (Metric: Accuracy $\uparrow$)
Credit $0.7341_{0.0085}$ $0.7431_{0.0102}$ $\underline{0.7441}_{0.0092}$ $\mathbf{0.7471}_{0.0092}$
Heloc $\underline{0.7107}_{0.0120}$ $0.7027_{0.0135}$ $\mathbf{0.7117}_{0.0145}$ $0.7077_{0.0113}$
Electricity $0.7762_{0.0070}$ $\mathbf{0.7962}_{0.0085}$ $\underline{0.7932}_{0.0075}$ $0.7862_{0.0087}$
Phoneme $0.7742_{0.0210}$ $\mathbf{0.8292}_{0.0245}$ $0.7942_{0.0207}$ $\underline{0.8242}_{0.0252}$
MagicTelescope $0.8444_{0.0088}$ $\underline{0.8479}_{0.0095}$ $0.8389_{0.0092}$ $\mathbf{0.8495}_{0.0079}$
Vehicle $\underline{0.8106}_{0.0340}$ $0.7576_{0.0380}$ $0.7676_{0.0362}$ $\mathbf{0.8176}_{0.0354}$
Oranges-vs-Grape $0.9105_{0.0150}$ $\underline{0.9302}_{0.0210}$ $0.9235_{0.0160}$ $\mathbf{0.9344}_{0.0240}$
Eye Movements $0.5699_{0.0135}$ $0.5819_{0.0155}$ $\underline{0.5899}_{0.0145}$ $\mathbf{0.5940}_{0.0164}$